OpenStudy (tatumlee):

5x-102=13 I know this is a basic equation but I'm very rusty on algebra can someone help me with this?

OpenStudy (photon336):

@tatumlee sure

OpenStudy (photon336):

step one, we must isolate x, we do this by adding 102, (which is the opposite of addition) to both sides. 5x-102 = 13 + 102 +102 5x = 115 5 5 divide both sides by 5 x = 23

OpenStudy (anonymous):

5x-102=13 first, add 102 both side 5x=115 then divide both side by 5 to gt the x by itself and there is the answer

OpenStudy (tatumlee):

And that works for most basic equations in algebra?

OpenStudy (photon336):

@tatumlee the idea behind these equations is that you must always isolate x, and whatever operation you do to one side, you must do to the other

OpenStudy (photon336):

@tatumlee if I gave you say kind of similar to what you had before 2x+8 = 20 how would you solve it?

OpenStudy (tatumlee):

I would add 20 to both sides?

OpenStudy (photon336):

@tatumlee we need to get x BY itself FIRST

OpenStudy (photon336):

So we would subtract 8 from both sides

OpenStudy (tatumlee):

How do you know what parts to add to both sides or subject or something

OpenStudy (photon336):

yeah that's a good question, you look at what you have in the equation

OpenStudy (anonymous):

subtract 8 both side and you'll get 2x=12 divide both side by 2

OpenStudy (tatumlee):

so x=6?

OpenStudy (anonymous):

when you look at the equation 2x+8=20, you want to get the 2x by itself in one side, so the only way you can do that is to subtract 8 to cancel out the 8's or make it zero. You see that positive8 - 8=0. The rule is that if you subtract 8 one side, you MUST do the same to the other side of the equation too. If the 8 was a negative (-8) , you would add 8 to make them zero. hope you understand it

OpenStudy (tatumlee):

I do thank you!

OpenStudy (anonymous):

pleasure